### v1

@@ 5.2 Plan Quality Comparisons
The same -point scoring system used to calculate rewards during DRL training was employed to quantify the plan quality of treatment plans generated for each patient’s new anatomy captured in the rpCT. The resulting plan quality scores are presented in Table 3. The score for each dosimetric metric of the generated plans are summarized in Appendix A. As shown, both DQN and PPO agents effectively adjusted the priorities of planning objectives, significantly improving the plan quality compared to the initial plans generated using the default priority set. Notably, for every case, both agents achieved plans that were comparable to or better than those manually created by a human planner. PPO-generated plans consistently achieved the highest quality scores across all cases (mean: ), outperforming both manual plans (, via paired t-test) and DQN-generated plans (). Furthermore, the PPO-generated plans exhibited lower score variability (, vs. for manual plans and for DQN-generated plans), indicating enhanced consistency and reliability in adapting to anatomical changes.
@@ 5.3 Dosimetric Analysis
The detailed dosimetric metrics of each structure for the manually generated plans and those generated by the DQN and PPO agents are presented in Table 4. Clinically acceptable target dose coverage was achieved in all the three plans for every case, with the coverage of primary CTV () no less than . In terms of dose homogeneity inside the primary CTV (CTV1), both RL agents reduced hot spots in three patients while maintaining excellent coverage. Most notably, in patient P4, DQN and PPO reduced the maximum dose within CTV1 by and , respectively, compared to the manual plan. PPO demonstrated superior OAR sparing in most patients, reducing mean dose to the larynx (LAR) ( (PPO) vs. (Manual)) and to the bilateral parotids (PARL and PARR) ( (PPO) vs. (Manual)), compared to manual replanning. The effectiveness of PPO was particularly evident in case P5, where it slightly compromised the dose coverage of the tertiary CTV (CTV3, ) to substantially reduce the dose to the right submandibular gland (SMGR), bringing its mean dose closer to the dose tolerance ( ) and resulting in a significantly higher plan score. In contrast, the manual and DQN-generated plans for this case substantially exceeded the dose tolerance of the SMGR, with mean doses of and , respectively. Figure 1 compares the DVH of the plans generated manually by a human planner against the automated plans generated using RL agents (DQN and PPO) for patients P4 and P5. DVH curves and dosimetric results for all cases are provided in Appendix A.
@@ 6 Discussion
Our study demonstrates that RL-based automated IMPT replanning, particularly using PPO, consistently generates treatment plans of higher quality for HNC patients experiencing anatomical changes during the treatment course, compared to those manually generated by a human planner. These improvements have several important implications for adaptive proton therapy workflows.
@@ 6.1 Clinical Significance
The marked improvement in plan quality scores across all patients indicates that RL-based automated replanning can standardize high-quality treatment planning. This may lead to reduced treatment-related toxicities while maintaining effective tumor control. For example, the reduction in parotid gland dose has been correlated with decreased risk of xerostomia, with studies suggesting that each reduction in mean dose translates to approximately reduction in xerostomia risk (Castelli et al., 2015; Chao et al., 2001). As shown in Table 4, patients P1 and P2 benefited from substantial parotid sparing (reductions of and , respectively), while patients P1 and P4 experienced notable laryngeal sparing (reductions of and ). Importantly, these dosimetric improvements were achieved without compromising target coverage, underscoring the ability of RL methods to address the geometric and dosimetric complexities in HNC treatment planning and balance complex trade-offs to achieve clinically acceptable treatment plans. The proposed RL-based automated replanning approach offers a promising solution for efficient offline adaptation and paves the way for online adaptive proton therapy.
@@ 6.2 RL-based Planning and Patient-specific Benefits
Both RL algorithms outperformed manual replanning, with PPO consistently delivering superior plans compared to DQN despite operating within the same discrete action space. This finding aligns with recent literature suggesting policy-based methods like PPO effectively navigate the complex reward landscapes common in radiotherapy planning (Wang and Chang, 2024). PPO’s strategy of performing small, incremental parameter updates helps prevent detrimental large policy shifts, a feature particularly well-suited to the radiotherapy optimization problem.
Analysis of patient-specific outcomes relative to anatomical features (Tables 2 and 3) underscores this advantage. PPO consistently delivered the highest plan scores, outperforming manual replans across the cohort irrespective of initial target volumes (rpCT CTV1 range: ) or the extent of their volumetric changes between pCT and rpCT (e.g., P2 CTV3: expansion). Notably, it achieved the highest absolute score () for the patient with the smallest CTV1/CTV2 (P4), while providing clinically meaningful gains over manual plans for large-volume cases (PPO plan score vs. manual for P3: ; P5: ). Overall, the results strongly suggest that DRL-based automation offers significant dosimetric advantages for IMPT replanning. By consistently optimizing plans towards high scores, the DRL framework holds considerable promise for improving plan quality for HNC patients undergoing significant anatomical changes.
@@ 6.3 Limitations and Future Directions
While our findings demonstrate the feasibility of deep reinforcement learning (DRL) for adaptive IMPT replanning, important limitations should be acknowledged alongside opportunities for future work. First, the retrospective evaluation was conducted on a limited cohort of five patients. However, this cohort is representative of challenging HNC cases with three CTVs of different prescription levels and common anatomical change patterns. Future work should apply our patient-specific approach to a larger, more diverse cohort of patients with varying CTV configurations, rigorously assessing the framework’s adaptability across different anatomical presentations. The demonstrated superiority of our patient-specific DRL agents over manual plans further motivates comprehensive evaluation in prospective studies.
@@ 7 Conclusion
R01DE033512 and R37CA272755.
Appendix A Complete Dosimetric Results

### v2

@@ 5.2 Plan Quality Comparisons
The same -point scoring system used to calculate rewards during DRL training was employed to quantify the plan quality of treatment plans generated for each patient’s new anatomy captured in the rpCT. To provide comprehensive benchmarking, we evaluated our patient-specific approach against both manual planning and population-based RL baselines. The population-based agents were trained on the initial planning CTs of the remaining patients in our cohort using identical architecture, hyperparameters, and training setup, consistent with conventional clinical practice where models are trained on population data.
[…]
Overall, patient-specific agents matched or exceeded manual plans in all cases, with PPO achieving the highest score in out of patients. Beyond manual benchmarks, patient-specific agents significantly outperformed population-based agents in of the comparisons (across all patients and algorithms), underscoring the clear advantages of personalized adaptation over generalized models, especially when handling unique anatomical changes.
@@ 5.3 Dosimetric Analysis
[…]
@@
5.4 Adaptation to Inter-fractional Changes
To further assess the robustness of our approaches in handling rapidly evolving anatomies, we extended our comparison to include the second replanning CT for cases in our evaluation cohort that underwent multiple replanning sessions during the treatment course (P1, P4, P6, P7, P8). As shown in Table 5, patient-specific RL demonstrated superior adaptation capabilities, achieving better performance than population-based baselines in out of metric comparisons (DQN: cases; PPO: cases). Patient-specific DQN, in particular, achieved a higher mean score of compared to for the population-based DQN baseline, while showing markedly lower variability. Although population-based approaches showed efficacy in handling certain cases with common anatomical patterns (e.g., P4 with DQN), the patient-specific approach consistently demonstrated better adaptation to the unique inter-fractional changes of individual patients.
Case ID | DQN-popn | DQN | PPO-popn | PPO |
P1 | 124.95 | 133.33 | 126.78 | 133.64 |
P4 | 146.23 | 143.13 | 145.11 | 145.88 |
P6 | 132.55 | 135.69 | 135.16 | 136.54 |
P7 | 115.27 | 140.57 | 135.21 | 141.11 |
P8 | 119.92 | 139.42 | 130.26 | 138.92 |
mean std |
@@ 6 Discussion
Our study demonstrates that patient-specific RL-based automated IMPT replanning, particularly using PPO, consistently generates treatment plans of higher quality for HNC patients experiencing anatomical changes during the treatment course, compared to both manually generated plans and population-based RL approaches. These improvements have several important implications for adaptive proton therapy workflows.
@@ 6.1 Clinical Significance
The marked improvement in plan quality scores across all patients indicates that patient-specific RL-based automated replanning can standardize high-quality treatment planning tailored to individual anatomical variations. This personalized approach may lead to reduced treatment-related toxicities while maintaining effective tumor control. For example, the reduction in parotid gland dose has been correlated with decreased risk of xerostomia, with studies suggesting that each reduction in mean dose translates to approximately reduction in xerostomia risk (Castelli et al., 2015; Chao et al., 2001). As shown in Table 4, patients P1 and P2 benefited from substantial parotid sparing (reductions of and , respectively), while patients P1 and P4 experienced notable laryngeal sparing (reductions of and ).
Importantly, these dosimetric improvements were achieved without compromising target coverage, underscoring the ability of patient-specific RL methods to address the geometric and dosimetric complexities unique to each patient’s HNC anatomy and balance complex trade-offs to achieve clinically acceptable treatment plans. The superior performance of patient-specific agents over population-based approaches ( comparisons in Table 3) demonstrates that individualized adaptation is crucial for handling the substantial inter-patient anatomical variability characteristic of HNC. The proposed patient-specific RL-based automated replanning approach offers a promising solution for efficient offline adaptation and paves the way for personalized online adaptive proton therapy.
@@ 6.2 RL-based Planning and Patient-specific Benefits
Patient-specific RL algorithms outperformed both manual replanning and population-based counterparts, with PPO consistently delivering superior plans compared to DQN. This finding aligns with recent literature suggesting policy-based methods like PPO effectively navigate the complex reward landscapes common in radiotherapy planning (Wang and Chang, 2024). PPO’s strategy of performing small, incremental parameter updates helps prevent detrimental large policy shifts, a feature particularly well-suited to patient-specific radiotherapy optimization where individual anatomical nuances require careful navigation.
[…]
These findings collectively demonstrate that patient-specific DRL delivers significant dosimetric advantages over population-averaged approaches by learning and adapting to individual patient characteristics. By consistently delivering high-quality plans tailored to each patient’s evolving anatomy, this framework holds considerable promise for improving adaptive radiotherapy outcomes in HNC treatment.
@@ 6.3 Limitations and Future Directions
While our findings demonstrate the feasibility of deep reinforcement learning (DRL) for adaptive IMPT replanning, important limitations should be acknowledged alongside opportunities for future work. First, the retrospective evaluation was conducted on a limited cohort of eight patients. However, this cohort is representative of challenging HNC cases with three CTVs of different prescription levels and common anatomical change patterns. Future work should apply our patient-specific approach to a larger, more diverse cohort of patients with varying CTV configurations, rigorously assessing the framework’s adaptability across different anatomical presentations. The demonstrated superiority of our patient-specific DRL agents over manual plans further motivates comprehensive evaluation in prospective studies.
@@ 7 Conclusion
R01DE033512 and R37CA272755. We thank Dr. David S. Yu, Dr. Mark McDonald, and Dr. Ralph Weichselbaum for providing valuable clinical perspectives that informed this study.
Appendix A Algorithm: Patient-Specific DRL for IMPT Replanning
Appendix B Cohort Demographics and Staging
Table 6 summarizes the demographic and clinical staging data for the eight patients included in this study. Patient ages range from to years. The cohort encompasses a range of disease stages, including stage II (), stage III (), stage IVa (), stage IVb (), and one unassigned case. Detailed TNM classifications are also reported, reflecting the heterogeneity of tumor and nodal involvement across the sample.
Case ID | Age | Overall Stage | TNM Stage |
P1 | 69 | IVb | T0 N3 M0 |
P2 | 73 | Unassigned | T4 & T2 N2 M0 |
P3 | 60 | II | T3 N2 M0 |
P4 | 55 | III | T4 N0 M0 |
P5 | 48 | II | T2 N2 M0 |
P6 | 72 | IVa | T4 N0 M0 |
P7 | 71 | III | T4 N0 M0 |
P8 | 59 | III | T4 N1 M0 |
Appendix C Priority Adjustments: Structure-Specific Modifications
All replanning agents modify the treatment plan optimization priorities through discrete adjustments to weight parameters. The complete mapping is detailed in Table 7, organized by structure type and adjustment magnitude. This discrete action space enables clinically meaningful trade-off adjustments while maintaining valid priority ranges. For critical OARs (brainstem, spinal cord), larger adjustments () are available to prioritize their protection, while other structures have more moderate adjustments. Target volumes include both positive and negative adjustments to allow for both improved coverage and compromises when necessary for OAR sparing.
Action Index | Structure | Parameter | Adjustment () |
Target Volume Adjustments | |||
0, 1 | CTV1 | ||
2, 3 | CTV2 | ||
4, 5 | CTV3 | ||
18, 19 | CTV2 | ||
20, 21 | CTV3 | ||
Organ-at-Risk Adjustments | |||
6 | Mandible (MAN) | ||
7 | Brainstem (BRS) | ||
8 | Spinal Cord (SC) | ||
9, 10 | Parotids (PARR, PARL) | ||
11 | Larynx (LAR) | ||
12 | Pharynx (PHY) | ||
13, 14 | Cochleas (COCHL, COCHR) | ||
15, 16 | Submand. Glands (SMGL, SMGR) | ||
17 | Esophagus (ESO) |
Appendix D Plan Quality Metrics
Plan quality was quantified using a comprehensive 150-point scoring system derived from standardized ProKnow criteria (Nelms et al., 2012) and institutional planning guidelines. This total score is the sum of individual component scores (detailed in Figure 2), which directly relate to the planning objectives outlined in Table 1. For reinforcement learning, the reward signal was defined as the change in this cumulative plan quality score between tuning steps.
Appendix E Network Architectures & Training Parameters
Both DRL agents process the DVHs using Convolutional Neural Networks (CNNs). This architectural choice is motivated by the ability of CNNs to learn meaningful spatial features and correlations from the structured DVH input, capturing complex dose-volume relationships across different anatomical structures. Common elements in both architectures include ReLU activation functions, batch normalization layers following convolutions, max-pooling for dimensionality reduction, and dropout layers for regularization. The DQN agent employs two convolutional layers followed by three fully connected layers to estimate Q-values for discrete actions. The DQN implementation features experience replay with prioritized sampling based on temporal difference errors, exponential epsilon-greedy decay, and a delayed target network updated every steps to stabilize learning. The PPO agent adopts a shared actor-critic architecture with a three-layer CNN backbone. The actor head outputs discrete priority adjustments via a categorical distribution, while the critic estimates state values. Training incorporates clipped policy updates (), generalized advantage estimation (GAE, ), value loss coefficient (), and entropy coefficient () to balance exploration and exploitation. Orthogonal initialization was applied to weights in the PPO network, with constant initialization for biases. Both algorithms use the Adam optimizer with learning rate , with PPO additionally employing gradient clipping (max norm ) to ensure training stability. Table 8 details the architectural specifications for both models. Table 9 highlights the key hyperparameters and optimization settings used in training.
Component | DQN | PPO |
Input Dimension | 15 × 100 (Structures × Dose Bins) | |
Input Channels | 1 | 1 |
Convolutional Layers | 2 | 3 |
Kernel Sizes | (3,5) | (3,5) |
Padding | (1,2) | (1,2) |
Pooling | 2×2 MaxPool | 2×2 MaxPool |
BatchNorm | Yes | Yes |
Dropout Rate | 0.2 | 0.2 |
Hidden Dim (CNN) | 32 → 64 | 32 → 64 → 64 |
Fully Connected | 256 → 128 → 22 | 512 → 22 (Actor) |
Output Activation | Linear | Softmax (Actor) |
Parameter | DQN | PPO | ||
Optimizer | Adam | Adam | ||
Learning Rate | 1e-5 | 1e-5 | ||
Minibatch Size | ||||
Discount Factor () | ||||
Loss Function | MSE |
(Eqn. 3) | (Eqn. 5) | |||
Experience Buffer | Prioritized Replay | On-policy Rollouts | ||
Buffer Retention | Long-term | Episodic | ||
Epsilon-Greedy Exploration | Yes (exponential decay) | No | ||
Advantage Estimation | – | GAE () | ||
Clipped Policy Update | – | |||
Value Clip | – | |||
Value Loss Coefficient () | – | |||
Entropy Coefficient () | – | |||
Gradient Clipping | Optional | (max norm) |
Appendix F Complete Dosimetric Results
[…]
@@ Appendix F Complete Dosimetric Results
Delineated | Dosimetric Endpoint | Manual | DQN | PPO | ||||
Structure | Metric | Score | Value | Score | Value | Score | Value | Score |
CTV1 | 15 | 98.00 | 15.00 | 98.01 | 15.00 | 98.00 | 15.00 | |
5 | 84.54 | 2.92 | 84.74 | 2.89 | 85.85 | 2.74 | ||
CTV2 | 15 | 98.69 | 15.00 | 98.41 | 15.00 | 98.51 | 15.00 | |
CTV3 | 15 | 98.46 | 15.00 | 98.15 | 15.00 | 98.39 | 15.00 | |
BRS | 21 | 1.11 | 21.00 | 1.03 | 21.00 | 1.23 | 21.00 | |
SC | 21 | 45.13 | 13.44 | 45.38 | 13.31 | 45.43 | 13.29 | |
MAN | 5 | 0.00 | 5.00 | 0.00 | 5.00 | 0.00 | 5.00 | |
LAR | 7 | 0.00 | 7.00 | 0.00 | 7.00 | 0.00 | 7.00 | |
PHY | 7 | 0.00 | 7.00 | 0.00 | 7.00 | 0.00 | 7.00 | |
PARL | 7 | 13.20 | 7.00 | 14.66 | 7.00 | 14.80 | 7.00 | |
PARR | 7 | 13.67 | 7.00 | 14.13 | 7.00 | 14.31 | 7.00 | |
COCHL | 5 | 0.00 | 5.00 | 0.00 | 5.00 | 0.00 | 5.00 | |
COCHR | 5 | 0.00 | 5.00 | 0.00 | 5.00 | 0.00 | 5.00 | |
SMGL | 5 | 36.19 | 4.53 | 34.94 | 5.00 | 34.97 | 5.00 | |
SMGR | 5 | 39.59 | 3.16 | 35.01 | 5.00 | 34.96 | 5.00 | |
ESO | 5 | 52.75 | 0.00 | 49.92 | 0.04 | 49.69 | 0.15 | |
Cumulative Total | 150 | – | 133.03 | – | 135.24 | – | 135.18 |
Delineated | Dosimetric Endpoint | Manual | DQN | PPO | ||||
Structure | Metric | Score | Value | Score | Value | Score | Value | Score |
CTV1 | 15 | 98.01 | 15.00 | 97.99 | 14.99 | 97.99 | 14.99 | |
5 | 78.64 | 3.77 | 82.44 | 3.22 | 84.30 | 2.96 | ||
CTV2 | 15 | 97.63 | 14.76 | 96.98 | 14.32 | 96.86 | 14.24 | |
CTV3 | 15 | 98.08 | 15.00 | 97.16 | 14.44 | 97.30 | 14.53 | |
BRS | 21 | 23.07 | 21.00 | 19.85 | 21.00 | 20.23 | 21.00 | |
SC | 21 | 34.99 | 18.50 | 30.29 | 20.85 | 30.27 | 20.87 | |
MAN | 5 | 0.03 | 5.00 | 0.07 | 5.00 | 0.21 | 5.00 | |
LAR | 7 | 42.36 | 6.05 | 41.64 | 6.34 | 40.61 | 6.76 | |
PHY | 7 | 50.47 | 6.81 | 48.31 | 7.00 | 48.50 | 7.00 | |
PARL | 7 | 35.68 | 0.00 | 25.86 | 7.00 | 25.61 | 7.00 | |
PARR | 7 | 23.26 | 7.00 | 23.10 | 7.00 | 25.32 | 7.00 | |
COCHL | 5 | 20.38 | 5.00 | 18.11 | 5.00 | 18.43 | 5.00 | |
COCHR | 5 | 7.54 | 5.00 | 5.31 | 5.00 | 5.60 | 5.00 | |
SMGL | 5 | 67.71 | 5.00 | 66.30 | 0.00 | 66.60 | 0.00 | |
SMGR | 5 | 35.91 | 4.63 | 35.35 | 4.86 | 34.84 | 5.00 | |
ESO | 5 | 14.48 | 5.00 | 12.24 | 5.00 | 12.65 | 5.00 | |
Cumulative Total | 150 | – | 132.53 | – | 141.03 | – | 141.35 |
Delineated | Dosimetric Endpoint | Manual | DQN | PPO | ||||
Structure | Metric | Score | Value | Score | Value | Score | Value | Score |
CTV1 | 15 | 97.99 | 14.99 | 97.99 | 14.99 | 98.02 | 15.00 | |
5 | 82.89 | 3.16 | 84.47 | 2.93 | 83.44 | 3.08 | ||
CTV2 | 15 | 97.86 | 14.91 | 98.06 | 15.00 | 97.99 | 14.99 | |
CTV3 | 15 | 97.77 | 14.85 | 97.51 | 14.68 | 97.51 | 14.68 | |
BRS | 21 | 15.29 | 21.00 | 15.94 | 21.00 | 15.74 | 21.00 | |
SC | 21 | 31.87 | 20.06 | 32.14 | 19.93 | 31.85 | 20.07 | |
MAN | 5 | 0.46 | 5.00 | 0.52 | 5.00 | 0.46 | 5.00 | |
LAR | 7 | 41.84 | 6.26 | 38.99 | 7.00 | 39.90 | 7.00 | |
PHY | 7 | 0.00 | 7.00 | 0.00 | 7.00 | 0.00 | 7.00 | |
PARL | 7 | 28.00 | 5.00 | 25.98 | 7.00 | 25.86 | 7.00 | |
PARR | 7 | 17.32 | 7.00 | 17.20 | 7.00 | 17.31 | 7.00 | |
COCHL | 5 | 4.15 | 5.00 | 4.40 | 5.00 | 4.37 | 5.00 | |
COCHR | 5 | 0.00 | 5.00 | 0.00 | 5.00 | 0.00 | 5.00 | |
SMGL | 5 | 69.03 | 0.00 | 68.98 | 0.00 | 68.77 | 0.00 | |
SMGR | 5 | 35.51 | 4.80 | 34.32 | 5.00 | 35.11 | 4.96 | |
ESO | 5 | 6.37 | 5.00 | 6.53 | 5.00 | 6.46 | 5.00 | |
Cumulative Total | 150 | – | 139.03 | – | 141.53 | – | 141.78 |
Structure | Metric | P6 | P7 | P8 | ||||||
M | Q | P | M | Q | P | M | Q | P | ||
CTV1 | 98.00 | 98.01 | 98.00 | 97.99 | 97.99 | 98.00 | 98.01 | 97.99 | 97.99 | |
84.54 | 84.74 | 85.85 | 82.89 | 84.47 | 83.44 | 78.64 | 82.44 | 84.30 | ||
CTV2 | 98.69 | 98.41 | 98.51 | 97.86 | 98.06 | 97.99 | 97.63 | 96.98 | 96.86 | |
CTV3 | 98.46 | 98.15 | 98.39 | 97.77 | 97.51 | 97.51 | 98.08 | 97.16 | 97.30 | |
BRS | 1.11 | 1.03 | 1.23 | 15.29 | 15.94 | 15.74 | 23.07 | 19.85 | 20.23 | |
SC | 45.13 | 45.38 | 45.43 | 31.87 | 32.14 | 31.85 | 34.99 | 30.29 | 30.27 | |
MAN | 0.00 | 0.00 | 0.00 | 0.46 | 0.52 | 0.46 | 0.03 | 0.07 | 0.21 | |
LAR | 0.00 | 0.00 | 0.00 | 41.84 | 38.99 | 39.90 | 42.36 | 41.64 | 40.61 | |
PHY | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 50.47 | 48.31 | 48.50 | |
PARL | 13.20 | 14.66 | 14.80 | 28.00 | 25.98 | 25.86 | 35.68 | 25.86 | 25.61 | |
PARR | 13.67 | 14.13 | 14.31 | 17.32 | 17.20 | 17.31 | 23.26 | 23.10 | 25.32 | |
COCHL | 0.00 | 0.00 | 0.00 | 4.15 | 4.40 | 4.37 | 20.38 | 18.11 | 18.43 | |
COCHR | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 7.54 | 5.31 | 5.60 | |
SMGL | 36.19 | 34.94 | 34.97 | 69.03 | 68.98 | 68.77 | 67.71 | 66.30 | 66.60 | |
SMGR | 39.59 | 35.01 | 34.96 | 35.51 | 34.32 | 35.11 | 35.91 | 35.35 | 34.84 | |
ESO | 52.57 | 49.92 | 49.69 | 6.37 | 6.53 | 6.46 | 14.48 | 12.24 | 12.65 | |
Plan Score (max=150): | 133.03 | 135.24 | 135.18 | 132.53 | 141.03 | 141.35 | 139.03 | 141.53 | 141.78 |